OpenStudy (anonymous):

In 1989, a movie ticket cost $3.00. In 1999, a movie ticket cost $5.50. Find the average rate of change of movie ticket prices in dollars per year. A. $0.20 B. $0.25 C. $4.00 D. $0.85

OpenStudy (anonymous):

new price-old price/newer year-older year

OpenStudy (anonymous):

5.5-3/1999-1989 = \[\frac{5.5-3}{1999-1989}\]

OpenStudy (anonymous):

\[\frac{2.5}{10}=.25\]